OpenStudy (anonymous):

The half-life of polonium is 139 days, but your sample will not be useful to you after 95% of the radioactive nuclei present on the day the sample arrives has disintegreated. For about how many days after the sample arrives will you be able to use the polonium? Formulas: Half-life = ln2/ k and y=yoe^-kt

OpenStudy (anonymous):

half life=ln2/k =139 k=ln2/139=0.005 approximately y=y0e^-kt, 5percent was left to you and usable y/yo)=5%=0.05=e^-kt ln(0.05)=lne^-kt ln(0.05)=-kt t=ln(0.05)/(-k) t=ln(0.05)/(-0.005) t=600 days approx ans....